# A Little Human Data Goes A Long Way

### v1

@@ Abstract
Faced with an expensive human annotation process, creators of NLP systems increasingly turn to synthetic data generation. While this method shows promise, the extent to which synthetic data can replace human annotation is poorly understood. We investigate the use of synthetic data in Fact Verification (FV) and Question Answering (QA) by studying the effects of incrementally replacing human generated data with synthetic points on eight diverse datasets. Strikingly, replacing up to 90% of the training data only marginally decreases performance, but replacing the final 10% leads to severe declines. We find that models trained on purely synthetic data can be reliably improved by including as few as 125 human generated data points. We show that matching the performance gain of just a little additional human data requires an order of magnitude more synthetic data, and we then estimate price ratios at which human annotation would be a more cost-effective solution. Our results suggest that even when human annotation at scale is infeasible, there is great value to having a small proportion of the dataset being human generated.
@@ 1 Introduction
From BERT (Devlin et al., 2019) to GPT-4 (Achiam et al., 2023), the explosive growth of language models (LMs) has been underpinned by exponential increases in the size of available training data. However, the more complex and specialized the task, the more expensive and challenging it is to collect human generated data at scale (Wang et al., 2021). Combined with growing concerns that LMs may soon exhaust the stock of publicly available training data (Villalobos et al., 2024), many turn to synthetic data generation, hoping to eliminate their reliance on human annotation.
@@
Modern LMs demonstrate the capability to solve myriad NLP tasks with minimal task specific data (Brown, 2020; Wei et al., 2022a, b; Ouyang et al., 2022; Ashok and Lipton, 2023), making them more powerful synthetic data generators. Leveraging this, synthetic data approaches have seen increased use in tasks (Tan et al., 2024) such as QA (Wu et al., 2021; Yu et al., 2024a), NLI (Meng et al., 2022; Ye et al., 2022; Hsieh et al., 2023), text classification (Ye et al., 2022; Gao et al., 2022; Yu et al., 2023, 2024b), instruction tuning (Li et al., 2024; Wei et al., 2023), evaluation (Dubois et al., 2024; Fu et al., 2023), and more (Tang et al., 2023; Kulkarni et al., 2024; Chan et al., 2024).
@@
The adoption has been particularly enthusiastic for tasks that require the model to ‘understand’ knowledge contained in an ‘evidence text’ e.g. FV (Tang et al., 2024), factual error correction (Ashok et al., 2023), NLI (Hosseini et al., 2024) and Context-grounded QA (Schimanski et al., 2024). Such tasks are of vital importance in Fake News Detection (Sharma et al., 2023), Retrieval Augmented Generation (Gao et al., 2023) and Dialogue Systems (Weston et al., 2015). Recent datasets (Wan et al., 2024; Wu et al., 2021; Liu et al., 2022; Ni et al., 2024; Muhlgay et al., 2023) and methods (Ye et al., 2022; Hsieh et al., 2023; Tang et al., 2024; Ni et al., 2024) exploit the plentiful availability of evidence texts (scientific journals, news articles, Wikipedia, books, etc.), using synthetic generation to avoid being bottlenecked by the expensive annotation procedure (Liu et al., 2022).
@@
Varying results from studies across ML tasks suggest that whether completely replacing human annotation with synthetic data shows promise (Fan et al., 2024; Hammoud et al., 2024) or leads to catastrophic failures (Bisbee et al., 2024; Guo et al., 2023) is highly task dependent (Li et al., 2023). In this work, we focus on FV and QA, investigating the trade-offs presented by the use of synthetic data generation in these fundamental tasks.
@@
We study eight diverse FV and QA datasets, using their ‘evidence texts’ to generate synthetic datasets. By holding the number of data points constant but increasing the percentage of the training data that is synthetic, we can compare the utility of synthetic data to the original human generated data points. Across multiple fine-tuning models, prompt models, and prompting strategies, we find that while increasing the proportion of synthetic data typically causes only minor degradations in model performance, a significant decline occurs at the extremes; i.e., when the percentage of synthetic data exceeds 90%. Focusing on the extremes, we show that purely synthetically trained FV and QA systems can be meaningfully improved by including as few as 125 human generated datapoints.
@@
Aside from the insight that synthetic data remains inferior to human annotation, our observations have actionable implications for researchers hoping to use synthetic data for FV and QA. Our results suggest that even when human annotation at scale is infeasible, there is great value to having a small proportion of the dataset being generated by humans.
@@
To help guide this choice, we quantify the performance-cost tradeoff between human and synthetic data. We find that matching the performance gain of just a little additional human data (only 200 data points) requires an order of magnitude more synthetic data points, empirically showing the per-data point price ratio at which human annotation is the more cost-effective solution. Finally, we conduct an analysis on the differing properties of synthetic vs. human data. Among other findings, we observe that synthetic generations are longer and more extractive from the evidence texts than their human-produced counterparts.
@@ 2 Synthetic Data Generation from Evidence Texts
We study a synthetic data generation pipeline representative of the methods used in the FV (Ni et al., 2024; He et al., 2023) and QA (Schimanski et al., 2024; Wan et al., 2024) literature. Specifically, we use Few-Shot In-Context Learning (Brown, 2020) to generate synthetic (claim, label) pairs from an input evidence text. The prompt model is given examples of (evidence text, claim, label) from the real training data, and is then queried with the evidence text we seek to generate data for (QA synthetic data is generated analogously, see additional details in Appendix B).
In total, we use 4 FV/NLI datasets: FEVER (Thorne et al., 2018), SciFact (Wadden et al., 2020), WANLI (Liu et al., 2022) and FACTIFY1.0 (Mishra et al., 2022), as well as 4 QA datasets: ROPES (Lin et al., 2019), CoQA (Reddy et al., 2019), QAConv (Wu et al., 2021) and FairyTaleQA (Xu et al., 2022). Together, the datasets span a variety of domains (science, news, social media, reasoning, conversation, fiction). More details, including a discussion on data leakage, can be found in Appendix C.
FV performance is measured by test accuracy, while QA is measured using Exact Match, String Inclusion, BLEU (Papineni et al., 2002), ROUGE-1 (Lin, 2004) and BERTScore (Zhang* et al., 2020). Evaluation is always conducted on the (human-generated) test split of each dataset.
We use GPT-3.5-Turbo (Brown, 2020) for prompting and LoRA (Hu et al., 2022) on Llama3-8B (Dubey et al., 2024) for fine-tuning. Implementation details are provided below (Appendix E). [^1] https://github.com/DhananjayAshok/LittleHumanData/
@@ 3 Can Synthetic Data Replace Humans?
Results: Across all datasets, using purely synthetic data typically leads to worse performance than the same amount of real data (Figure 1). This confirms that despite advances in synthetic generation, human annotation yields more useful data.
The performance decline is not uniform as we increase the synthetic proportion. On all datasets, there is only a minor degradation up until 90% replacement, after which the performance drops considerably. We zoom in on the 90%-100% interval, fixing the amount of training data at $n=5000$ (500 for SciFact) and training on datasets with 95%, 97.5% and 100% synthetic data (Figure 2). Surprisingly, the results show that across all datasets (and 3 random seeds), there is a significant difference between the performance of models on 97.5% and 100% synthetic data. These trends hold robustly over choice of fine-tuning model (Mistral7B), prompt model (GPT4), prompting strategy (Chain-of-Thought) and data scale (Appendix A). The same trend is seen in all cases; when using synthetic training sets, the addition of just 125 (2.5% of 5000) human generated datapoints reliably improved the performance of FV and QA models.
@@ 4 When Should We Use Human Data?
Having observed the disproportionate value added by human data, we ask what the relative cost between human and synthetic data generation must be for us to prefer one over the other. We fine tune models on purely synthetic datasets of varying sizes, and establish the synthetic baseline by fitting a curve of the form $y=a_{0}+a_{1}\log(x)$ where $x$ is the size of the synthetic dataset and $y$ is the performance. We then take the synthetic training sets with {1000, 2000 …} points and observe the performance ($y^{*}$) when we add 200 human data points. $\exp({\frac{y^{*}-a_{0}}{a1}})$ is then the size of the purely synthetic dataset that achieves equivalent performance.
Results: Across all datasets, adding 200 human data points is comparable to adding at least an order of magnitude (often multiple orders of magnitude) more synthetic data points. On WANLI (Figure 3), more than $17000$ additional synthetic points are needed to achieve the performance gains of $200$ human points. If the price of a synthetic point for WANLI exceeds $73$ times the price of a human generated point, then an incremental amount of human annotation would be the more cost-effective solution. In the extreme case, the equation learned on FairyTaleQA suggests that it takes $52e$ 5 additional synthetic points to match the performance gain of 200 additional real data points. Rather than interpret these numbers literally, we take them to suggest that human data could have unique value in some settings, enabling performance levels that are impossible with purely synthetic datasets. See below (Appendix A) for more results and details.
@@ 5 Discussion
A deeper analysis of the difference between real and synthetic data exposes interesting trends. Our analysis shows that synthetic data generation produced claims of comparable length to the real datasets, however synthetic questions and answers tended to be longer than human-generated counterparts for all QA datasets (Figure 4). We find that synthetic generations have a higher n-gram overlap with the evidence sentences. This suggests that synthetic data generation produces data points that are more directly taken from the evidence texts, while humans are more likely to employ rephrasing or different vocabulary than the evidence texts. Surprisingly, we find that synthetic data generation chooses more diverse sources for the question and answer content, with human annotation overwhelmingly more likely to create questions whose answers lie in the start of the evidence texts. We include a detailed discussion below (Appendix D).
@@ 6 Related Work
The replacement of human annotation with synthetic data is extensively studied in the pretraining stage of LMs, where results consistently show (Shumailov et al., 2023; Seddik et al., 2024; Guo et al., 2023; Briesch et al., 2023) catastrophic forgetting, mode collapse, and performance deterioration.
While our results show that purely synthetic datasets are relatively worse than mixed datasets, relying only on synthetic data still achieves reasonable performance across all tasks. This suggests that the usage of exclusively synthetic data poses fewer risks in a setting where generations are still grounded in diverse natural ‘evidence texts.’
@@
There is limited work on understanding whether synthetic data can replace human annotation in a task-specific setting for the language domain. Li et al. (2023) categorize text classification tasks by subjectivity, showing that synthetic data is less useful when tasks are more subjective. This draws them to focus on different tasks (sentiment classification, relation extraction and spam detection), and they do not study using a mixture of real and synthetic data.  Bisbee et al. (2024) demonstrate that replacing political survey respondents with LMs produces unreliable results, while  Ahmed et al. (2024) find that there are specific subtasks in the annotation of software engineering artefacts where synthetic data approaches human performance.  Chen et al. (2024) show that instruction-following capabilities are diminished when using data from GPT-4, and present a machine unlearning approach to mitigate this. The diversity of results when evaluating the impact of using purely synthetic data confirms that the feasibility of replacing human annotation with synthetic data is highly task dependant. This work deepens our understanding of the problem by being the first to study whether synthetic data can replace human annotation on the fundamental tasks of fact verification, question answering, and natural language inference.
@@ 8 Limitations
All datasets studied in this work are in the English language, hence limiting our scope of investigation. We hope future work can explore whether similar claims can be made regarding the impact of replacing human annotation with synthetic data across different languages. We also have a limited ability to control for dataset leakage, with only one dataset from each of the tasks that is surely not leaked to GPT-3.5 (and, even these two datasets may have been seen by GPT-4). This can potentially bias the results in favor of synthetic data. Due to the scarcity of suitable available datasets (i.e., ones that have not been exposed to the prompt models) we are prevented from studying the problem more rigorously. Another limitation is that while we are able to identify clear differences between synthetic vs. real data distributions, our analysis of the errors made by models trained on 0% vs. 100% synthetic data failed to yield any generalizable insights that could inform modelling approaches. A more fine-grained study of the effect of using synthetic data on the behaviour of the downstream model is hence left as a subject of future research. This work focuses on the value of human-labelled data, but due to the nature of benchmarks, the human-labelled test data is drawn from the same distribution as our treatment data. Future work should verify that these results hold under novel human-generated examples, which is beyond the scope of this investigation.

### v2

@@ Abstract
Faced with an expensive human annotation process, creators of NLP systems increasingly turn to synthetic data generation. While this method shows promise, the extent to which synthetic data can replace human annotation is poorly understood. We investigate the use of synthetic data in Fact Verification (FV) and Evidence-based Question Answering (QA) by incrementally replacing human-generated data with synthetic points on eight diverse datasets. Strikingly, replacing up to 90% of the training data only marginally decreases performance, but replacing the final 10% leads to severe declines. We find that models trained on purely synthetic data can be improved by including as few as 125 human generated data points. We show that matching the performance gain of a little human data requires an order of magnitude more synthetic data, and then estimate price ratios at which human annotation would be a more cost-effective solution. Our results suggest that even when human annotation at scale is infeasible, there is great value to having a small proportion of the dataset being human-generated.
Figure 1: Change in model performance as the proportion of synthetic points in the training data is increased. Across datasets, the performance decrease when moving the synthetic proportion from 0 to 0.90 is often less than that of moving from 0.9 to purely synthetic data.
@@ 1 Introduction
From BERT (Devlin et al., 2019) to GPT-4 (Achiam et al., 2023), the explosive growth of language models (LMs) has been underpinned by exponential increases in the size of available training data. However, the more complex and specialized the task, the more expensive and challenging it is to collect human-generated data at scale (Wang et al., 2021). Combined with growing concerns that LMs may soon exhaust the stock of publicly available training data (Villalobos et al., 2024), many turn to synthetic data generation, hoping to eliminate their reliance on human annotation.
@@
Modern LMs demonstrate the capability to solve myriad NLP tasks with minimal task specific data (Brown et al., 2020; Wei et al., 2022), making them more powerful synthetic data generators. Leveraging this, synthetic data approaches have seen increased use in tasks (Tan et al., 2024) such as QA (Wu et al., 2022), natural language inference (NLI) (Meng et al., 2022), text classification (Ye et al., 2022), instruction tuning (Li et al., 2024), evaluation (Dubois et al., 2024), and more (Tang et al., 2023).
Figure 2: Model performance as the synthetic proportion of the training data varies from 0.95 to 1. Having just 2.5% of the training dataset being human-generated boosts performance.
@@
The adoption has been particularly enthusiastic for tasks that require the model to ‘understand’ knowledge contained in an ‘evidence text’ e.g., FV (Tang et al., 2024), factual error correction (Ashok et al., 2023), NLI (Hosseini et al., 2024) and evidence-based QA (Schimanski et al., 2024). Such tasks are of vital importance in fake news detection (Sharma et al., 2023), retrieval-augmented generation (Gao et al., 2023) and dialogue systems (Weston et al., 2015). Recent datasets (Wu et al., 2022) and methods (Ye et al., 2022) exploit plentiful evidence texts (scientific journals, news articles, books, etc.), using synthetic generation to avoid being bottlenecked by the expensive annotation procedure (Liu et al., 2022).
@@
Varying results across ML tasks suggest that whether completely replacing humans with synthetic data shows promise (Fan et al., 2024; Hammoud et al., 2024) or leads to failures (Bisbee et al., 2024; Guo et al., 2024) is task dependent (Li et al., 2023). In this work, we focus on FV and Evidence-based QA, performing the first investigation into the trade-offs presented by the use of synthetic data generation in these fundamental tasks.
@@
We study eight diverse FV and QA datasets, using their ‘evidence texts’ to generate synthetic datasets. By holding the number of data points constant but increasing the percentage of the training data that is synthetic, we can compare the utility of synthetic data to the original human generated data points. Across multiple models, prompt models, and prompting strategies, we find (Figure 1) that while increasing the proportion of synthetic data typically causes only minor degradations in model performance, a significant decline occurs at the extremes; i.e., when the percentage of synthetic data exceeds 90%. Focusing on the extremes, we show that purely synthetically trained FV and QA systems can be meaningfully improved by including as few as 125 human-generated datapoints.
@@
Our observations have actionable implications for researchers hoping to use synthetic data for FV and QA. The results (Figure 2, Figure 4) suggest that even when human annotation at scale is infeasible, there is great value to having a small proportion of the dataset being generated by humans.
@@
To help guide this choice, we quantify the performance-cost tradeoff between human and synthetic data. We find (Figure 4) that matching the performance gain of just a little additional human data (only 200 data points) requires an order of magnitude more synthetic data points, empirically showing the per-data point price ratio at which human annotation is the more cost-effective solution. Finally, we conduct an analysis on the differing properties of synthetic vs. human data. Among other findings, we see that synthetic generations are longer and more extractive from the evidence texts than their human-produced counterparts.
@@ 2 Synthetic Data Generation from Evidence Texts
We study a synthetic generation pipeline representative of the methods used in the FV (Ni et al., 2024; He et al., 2023) and QA (Schimanski et al., 2024; Wan et al., 2024) literature. Using Few-Shot In-Context Learning (Brown et al., 2020), we generate synthetic (claim, label) pairs from an input evidence text. The prompt model is given examples of (evidence text, claim, label) from the human training data, and is then queried with the evidence text we seek to generate data for. QA synthetic data is generated analogously, see details in Appendix B.
In total, we use four FV/NLI datasets: FEVER (Thorne et al., 2018), SciFact (Wadden et al., 2020), WANLI (Liu et al., 2022) and FACTIFY1.0 (Mishra et al., 2022), as well as four QA datasets: ROPES (Lin et al., 2019), CoQA (Reddy et al., 2019), QAConv (Wu et al., 2022) and FairyTaleQA (Xu et al., 2022). Together, the datasets span a variety of domains (science, news, social media, reasoning, conversation, fiction). We confirm that the generations are of high quality by verifying that the diversity of the synthetic data is comparable to the human generated samples. For more details, including a discussion on data leakage, see Appendix C.
FV performance is measured by test accuracy, while QA is measured using BLEU (Papineni et al., 2002); we show robustness to choice of metric in Appendix A. Evaluation is always conducted on the (human-generated) test split of each dataset.
We use GPT-3.5-Turbo (Brown et al., 2020) for prompting and LoRA (Hu et al., 2022) on Llama3-8B (Dubey et al., 2024) for fine-tuning. Implementation details are provided in Appendix E, and our code is publicly available  [^1] https://github.com/DhananjayAshok/LittleHumanData/
Figure 3: Change in accuracy when the test set (shown in key) is not seen during training, and the training set is a mixture of other FV datasets. Increasing the synthetic proportion of the dataset leads to performance declines even in the OOD setting, showing that human data offers genuine performance increases.
@@ 3 Can Synthetic Data Replace Humans?
Figure 4: On the WANLI dataset, adding 200 real data points is as effective as adding an order of magnitude more synthetic data points.
Results: Across all datasets, using purely synthetic data leads to worse performance than the same amount of human data (Figure 1). We consider the possibility that this result could be caused by a spurious correlation between the human training and testing splits (e.g., annotation artifacts that are correlated with the label but not fundamental to the task). We conduct an out-of-distribution experiment, using different datasets for training and testing (e.g., training on FEVER + SciFACT and testing on WANLI). Increasing the synthetic proportion leads to performance declines even in the OOD setting (Figure 3), showing that human data offers genuine performance increases, and the results cannot be explained by a spurious correlation between the human test and human training samples (further discussion in Appendix A).
The performance decline is not uniform as we increase the synthetic proportion. On almost all datasets, there is only a minor degradation up until 90% replacement, after which the performance drops considerably. We zoom in on the 90%-100% interval, fixing the amount of training data at $n=5000$ (500 for SciFact) and training on datasets with 95%, 97.5% and 100% synthetic data (Figure 2). Surprisingly, the results show that there is a significant difference between the performance of models on 97.5% and 100% synthetic data; the addition of just 125 (2.5% of 5000) human generated datapoints reliably improves the performance of synthetically trained FV and QA models. These trends hold robustly over different languages (Arabic, Georgian, Indonesian), choice of fine-tuning model (Mistral, MPT), prompt model (GPT4 and Claude-3.5-Sonnet), prompting strategy (Chain-of-Thought), model size and dataset size (Appendix A).
@@ 4 When Should We Use Human Data?
Dataset Mean Median 25th Percentile 75th Percentile
WANLI 17,671 16,905 9,711 22,931
ROPES 17,333 6,006 3,623 21,944
FairyTaleQA 281,951 36,901 15,129 813,135
FEVER 1,155 237 -1,400 7,073
Table 1: Additional synthetic data points needed to match the performance gain of 200 human data points. High values for FairyTaleQA suggest that human-generated data may unlock performance that purely synthetic data cannot achieve. Negative values for FEVER are due to a saturation of the performance gains, however, human data points reach the saturation point much faster (Appendix A)
Having observed the disproportionate value added by human data, we ask what the relative cost between human and synthetic data generation must be for us to prefer one over the other. We fine-tune models on purely synthetic datasets of varying sizes, and establish the synthetic baseline by fitting a curve of the form $y=a_{0}+a_{1}\log(x)$ where $x$ is the size of the synthetic dataset and $y$ is the performance. We then take the synthetic training sets with {1000, 2000 …} points and observe the performance ($y^{*}$) when we add 200 human data points. $\exp({\frac{y^{*}-a_{0}}{a1}})$ is then the size of the purely synthetic dataset that achieves equivalent performance.
Dataset Synthetic Human
FEVER 5.50 20.27
WANLI 0.23 1.22
SCIFACT 0.00 9.93
FACTIFY 0.27 14.93
NarrativeQA 3.85 1.42
CoQA 0.54 5.49
FairyTaleQA 2.26 0.18
ROPES 2.40 1.35
Table 2: Percentage of duplicated claims/questions for synthetic v.s. human data. Rates are comparable across datasets, but for fact verification datasets, synthetic datasets have fewer duplicates.
Results: Across all datasets, adding 200 human data points is usually comparable to adding at least an order of magnitude (often multiple orders of magnitude) more synthetic data points. On WANLI (Figure 4), more than $17,000$ additional synthetic points are needed to achieve the performance gains of $200$ human points. If the price of a synthetic point for WANLI exceeds $73$ times the price of a human generated point, then an incremental amount of human annotation would be a more cost-effective way to achieve the same increase in accuracy. In the extreme case, the equation learned on FairyTaleQA suggests that it takes $52e$ 5 additional synthetic points to match the performance gain of 200 additional human data points. Rather than interpret these numbers literally, we take them to suggest that human data could have unique value in some settings, enabling performance levels that are impossible with purely synthetic datasets. See Appendix A for more results and details.
@@ 5 Discussion
The synthetic generations are as diverse as human data (Table 2), with comparable duplicate rates on QA datasets, and markedly fewer duplicates on most FV datasets. This is evidence that our synthetic generations are of good quality, however, even on the datasets where the synthetic data is significantly more diverse, the synthetic data does not perform as well. This suggests that diversity is an insufficient measure of quality when evaluating how good the generated data is. Our analysis shows that synthetic data generation produces claims of comparable length to the human datasets, however synthetic questions and answers tend to be longer than human-generated counterparts for all QA datasets (Figure 5). We find that synthetic generations have a higher n-gram overlap with the evidence sentences. This suggests that synthetic data generation produces data points that are more directly taken from the evidence texts, while humans are more likely to employ rephrasing or different vocabulary than the evidence texts. Surprisingly, we find that synthetic data generation chooses more varied parts of the input text as sources for the question and answer content, with human annotation overwhelmingly more likely to create questions whose answers lie at the start of the evidence texts. We include a detailed discussion in Appendix D.
@@ 6 Related Work
The replacement of human annotation with synthetic data is extensively studied in the pretraining stage of LMs, where results consistently show (Shumailov et al., 2023; Seddik et al., 2024; Guo et al., 2024; Briesch et al., 2023) catastrophic forgetting, mode collapse, and performance deterioration.
In our setting, relying only on synthetic data still achieves reasonable performance across all tasks. This suggests that the usage of exclusively synthetic data poses fewer risks when generations are grounded in diverse, natural ‘evidence texts.’
@@
There is limited work on understanding whether synthetic data can replace human annotation in a task-specific setting for the language domain. Li et al. (2023) categorize text classification tasks by subjectivity, showing that synthetic data is less useful when tasks are more subjective. This draws them to focus on different tasks (sentiment classification, relation extraction and spam detection), and they do not study using a mixture of real and synthetic data.  Bisbee et al. (2024) demonstrate that replacing political survey respondents with LMs produces unreliable results, while  Ahmed et al. (2024) find that there are specific software engineering subtasks where synthetic data approaches human performance.  Chen et al. (2024) show that instruction-following capabilities are diminished when using synthetic data and present a machine unlearning approach to mitigate this. The diversity of results when evaluating the impact of using purely synthetic data confirms that the feasibility of replacing human annotation with synthetic data is highly task dependant. This work deepens our understanding of the problem by being the first to study whether synthetic data can replace human annotation on the fundamental tasks of fact verification and evidence-based question answering.
@@ 8 Limitations
While we include results on multilingual Fact Verification datasets, the primary focus of our work is limited to the English language. Additionally, our results on multilingual datasets suggest that while similar claims can be made regarding the impact of replacing human annotation with synthetic data across different languages, the amount of human data needed to observe a meaningful performance increase may vary across languages. We also have a limited ability to control for dataset leakage, with only one dataset from each of the tasks that is surely not leaked to GPT-3.5 (and, even these two datasets may have been seen by GPT-4). This can potentially bias the results in favor of synthetic data. Due to the scarcity of suitable available datasets (i.e., ones that have not been exposed to the prompt models) we are prevented from studying the problem more rigorously. Another limitation is that while we are able to identify clear differences between synthetic vs. real data distributions, our analysis of the errors made by models trained on 0% vs. 100% synthetic data failed to yield any generalizable insights that could inform modelling approaches. A more fine-grained study of the effect of using synthetic data on the behaviour of the downstream model is hence left as a subject of future research.